OpenStudy (anonymous):

Prove the Identity: 1+sinx-cos^2x/1+sinx=sinx

OpenStudy (anonymous):

\[\frac{ 1+sinx+\sin^2x-1 }{sinx+1 } \] one cancel with minus one , take sin(x) as a common factor and see what happens

OpenStudy (anonymous):

\[\frac{1+\sin x-\cos^2x}{1+\sin x}=\sin x\] for those wondering

OpenStudy (anonymous):

...........................

OpenStudy (anonymous):

this will help you simplify it: \[\cos2x = \cos^2x-\sin^2x\]

OpenStudy (anonymous):

Ummmmmmmm I don't understand what you are getting at. I just need to get the left to equal the right........

OpenStudy (anonymous):

didn't you get my point ?

OpenStudy (anonymous):

No....I don't understand.

OpenStudy (anonymous):

you have sin^x+cos^x=1 ,so cos^x=1-sin^x , now we sub in the LHS\[\frac{ 1+sinx+\sin^2x-1 }{ sinx+1 } \rightarrow \frac{ sinx (sinx+1) }{ (sinx+1) }=sinx\]

OpenStudy (anonymous):

Nevermind.................

OpenStudy (anonymous):

just tell me what's the problem :(

OpenStudy (anonymous):

You just didn't explain it clear enough.........I got my question answered though, thanks.

OpenStudy (anonymous):

I'm sorry I've just sen it ... good luck

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Dear awesomelady2, you've go to know the rules about the trigonometric functions, the proof is clear enough, I would recommend you to see the geometrical explanations of those proofs.

OpenStudy (anonymous):

I'm just really stressed, that's all. I have a /GIANT/ test tomorrow